OpenStudy (anonymous):

Solve the system using matrices. 3x + 4y = 18 3x + 4y = 18 A. (-2, 6) B. (-2, -6) C. (6, -2) D. No solution

jimthompson5910 (jim_thompson5910):

you posted the same equation twice

OpenStudy (anonymous):

I didn't notice

jimthompson5910 (jim_thompson5910):

is that a typo?

jimthompson5910 (jim_thompson5910):

or does it really list 3x + 4y = 18 twice?

OpenStudy (anonymous):

no it's listed once

jimthompson5910 (jim_thompson5910):

what is the other equation

OpenStudy (anonymous):

This is the problem. Solve the system using matrices. 3x + 4y = 18 3x + 4y = 18 A. (-2, 6) B. (-2, -6) C. (6, -2) D. No solution

jimthompson5910 (jim_thompson5910):

I know, but what are the two equations

OpenStudy (anonymous):

I have no Idea I don't seem to understand the question

jimthompson5910 (jim_thompson5910):

one of the equations is 3x+4y = 18 what is the other one?

jimthompson5910 (jim_thompson5910):

if both equations are 3x + 4y = 18, then you'd get infinitely many solutions

OpenStudy (anonymous):

It is like this in the page 3x + 4y = 18 3x + 4y = 18

jimthompson5910 (jim_thompson5910):

then there's a typo

jimthompson5910 (jim_thompson5910):

I would bring it up with the teacher and see what s/he says

OpenStudy (anonymous):

It's no solution

jimthompson5910 (jim_thompson5910):

no actually quite the opposite: there are infinitely many solutions that's if the system is 3x + 4y = 18 3x + 4y = 18

OpenStudy (anonymous):

ok

jimthompson5910 (jim_thompson5910):

so that's why I think it's a typo and the second equation should be something else

OpenStudy (anonymous):

Do you know this one Solve the system of equations by the addition method. If the system does not have a single ordered pair as a solution, state whether the system is inconsistent or dependent. -4x - 3y = 6 -20x - 15y = 30 A. No solution, inconsistent B. (-24, -18) C. (-18, -24) D. Infinitely many solutions, dependent

jimthompson5910 (jim_thompson5910):

notice how if you multiplied everything in equation 1 by 5, you get equation 2 this means that the two equations are effectively the same

jimthompson5910 (jim_thompson5910):

so this means the system is dependent and there are infinitely many solutions

OpenStudy (anonymous):

Ok thanks :)

jimthompson5910 (jim_thompson5910):

yw

OpenStudy (anonymous):

How do you work this problem out am not good in math at all but I have to pass this class and I have been in it for 1 year and couple of months

jimthompson5910 (jim_thompson5910):

just keep practicing and you'll get better

jimthompson5910 (jim_thompson5910):

and try to translate everything to something you understand

jimthompson5910 (jim_thompson5910):

math is like a language: you have to practice it everyday and you have to keep using it also you have to translate things to something you're already comfortable with and build from there

OpenStudy (anonymous):

am almost done wit this class this is the only one am missing to finish thank God

jimthompson5910 (jim_thompson5910):

well even though you'll be done, you should still keep practicing because math will be used in the workplace (most likely)

OpenStudy (anonymous):

since it's an online class it's hard for me to understand

jimthompson5910 (jim_thompson5910):

yeah online class is tough because you don't have a teacher to ask questions

OpenStudy (anonymous):

the books online don't help me understand how to get the answer

OpenStudy (anonymous):

yeap that's true

jimthompson5910 (jim_thompson5910):

yeah that's another thing that's usually cryptic: textbooks

OpenStudy (anonymous):

yeah

OpenStudy (anonymous):

I got this one wrong I put B can u help me? Solve the equation for y. 7x - 3y = 4 A. y = x + B. y = 7x - 4 C. y = x + D. y = x -

jimthompson5910 (jim_thompson5910):

ok one sec

OpenStudy (anonymous):

And I got this one wrong too Solve the equation for y. 7x - 3y = 4 A. y = x + B. y = 7x - 4 C. y = x + D. y = x -

jimthompson5910 (jim_thompson5910):

nothing pasted properly, but that's ok

jimthompson5910 (jim_thompson5910):

\[\large 7x - 3y = 4\] \[\large -3y = 4-7x\] \[\large -3y = -7x + 4\] \[\large y = \frac{-7x + 4}{-3}\] \[\large y = \frac{-7x}{-3} + \frac{4}{-3}\] \[\large y = \frac{7}{3}x - \frac{4}{3}\]

OpenStudy (anonymous):

So the answer is D?

OpenStudy (anonymous):

I mean C?

jimthompson5910 (jim_thompson5910):

I can't tell because like I said, things didn't paste properly

jimthompson5910 (jim_thompson5910):

look at what you pasted to see what I'm talking about

OpenStudy (anonymous):

Oh ok

OpenStudy (anonymous):

A. y= x+ B. y= 7x -4 C. y=x + D. y= x -

OpenStudy (anonymous):

That's the answers to choose I had choose B n got it wrong

jimthompson5910 (jim_thompson5910):

again, it's not pasting properly post a screenshot instead

OpenStudy (anonymous):

how do I do a screenshot?

jimthompson5910 (jim_thompson5910):

hit printscreen open up MS Paint paste it in save it, then post it

jimthompson5910 (jim_thompson5910):

printscreen is the key around the insert, home, page up keys

OpenStudy (anonymous):

Ok

OpenStudy (anonymous):

Cant find it I guess cause it's a online class

jimthompson5910 (jim_thompson5910):

it's a key on your keyboard

OpenStudy (anonymous):

OK found it

OpenStudy (anonymous):

so I high light the question n then press printscreen?

jimthompson5910 (jim_thompson5910):

no just hit printscreen then you open up MS Paint and paste

OpenStudy (anonymous):

ok

jimthompson5910 (jim_thompson5910):

printscreen just copies the entire screen that you see and it shares it with whoever

OpenStudy (anonymous):

were is MS paint?

OpenStudy (anonymous):

found it

OpenStudy (anonymous):

my laptop says paint alone

jimthompson5910 (jim_thompson5910):

ah yeah paint works

jimthompson5910 (jim_thompson5910):

hmm very weird, it's not showing up there either

jimthompson5910 (jim_thompson5910):

still the same thing

jimthompson5910 (jim_thompson5910):

you'll have to contact a teacher about this it looks like there are some serious flaws with it

jimthompson5910 (jim_thompson5910):

the answer choices aren't even fully written out

OpenStudy (anonymous):

ok

jimthompson5910 (jim_thompson5910):

seems like its common for typos (really bad ones) to show up on this system

OpenStudy (anonymous):

thank for ur help

jimthompson5910 (jim_thompson5910):

sure thing

OpenStudy (anonymous):

can u help me with this one? Decide whether the ordered pair is a solution of the given system. (-1, 6) x + y = 7 x - y = -5 A. Yes B. No

jimthompson5910 (jim_thompson5910):

x + y = 7 -1 + 6 = 7 ... plug in x = -1 and y = 6 5 = 7 since that last equation is FALSE, this means that (-1,6) is NOT a solution to x+y = 7

jimthompson5910 (jim_thompson5910):

so (-1,6) is NOT a solution to the entire system

OpenStudy (anonymous):

thanks

OpenStudy (anonymous):

Perform the indicated operation.